# AlphaApollo: Orchestrating Foundation Models and Professional Tools into a Self-Evolving System for Deep Agentic Reasoning

### v1

@@ 4.3 Test-time Iteration
Most early test-time scaling methods utilize parallel iteration, where the model generates multiple independent reasoning trajectories. The final answer is then selected through various aggregation mechanisms. Self-Consistency (Wang et al., 2022) employs simple voting, Semantic Self-Consistency (Knappe et al., 2024) leverages semantic similarity, CISC (Taubenfeld et al., 2025) utilizes confidence metrics, and MCR (Yoran et al., 2023) implements model-decided final answers. Additionally, DIVSE (Naik et al., 2023) enhances the diversity of the reasoning trajectory by reformulating the original question to encourage exploration of a broader solution space. While this strategy enhances FM reasoning by allocating more computational budget to the reasoning process (Snell et al., 2024), it is fundamentally limited by the lack of interaction among different reasoning trajectories. This isolation limits the model in leveraging from previous attempts or refining its approach based on earlier outputs. The independence between trajectories constrains the potential for iterative improvement and fails to capitalize on insights that could emerge from comparing or combining intermediate reasoning steps (Qi et al., 2025).
@@
Sequential iteration strategies leverage prior reasoning processes and outcomes to identify limitations and iteratively refine subsequent generations. Notably, Self-Refine methods (Madaan et al., 2023; Qu et al., 2024; Chen et al., 2023c) prompt FMs to revise initial outputs, generating improved solutions. Similarly, Muennighoff et al. (2025) employs a deliberation mechanism, replacing the end-of-sequence token with the keyword ‘wait’ to enable continuous reasoning post-answer generation. OPRO (Yang et al., 2024b) iteratively optimizes prompts based on prior outcomes to yield superior solutions. While these approaches enhance reasoning, their reliance on self-correction without external supervision can lead to unreliable outcomes (Huang et al., 2024). In contrast, Reflexion (Shinn et al., 2023), RCI (Kim et al., 2023), and Refiner (Paul et al., 2023) incorporate external feedback from environments, code executors, and critic models, respectively, providing robust supervision and improving performance. Furthermore, methods like Alpha-Evolve (Novikov et al., 2025) and TextGrad (Yuksekgonul et al., 2025) extend sequential iteration to scientific domains, demonstrating FMs’ potential in tackling complex challenges, such as code optimization and molecular synthesis.
@@
Mixed iteration strategies integrate parallel and sequential iteration to combine the exploratory breadth of independent trajectories with the refinement depth of conditioned generations. This hybrid approach optimizes test-time compute allocation, often surpassing pure parallel or sequential methods on complex reasoning tasks by balancing exploration (through parallel iteration) and exploitation (via sequential refinement). A prominent example is Monte Carlo Tree Search (MCTS) (Silver et al., 2016, 2017), which employs multi-round node expansion through parallel iteration and outcome simulation via sequential iteration. Recent advancements apply MCTS to enhance FM reasoning by treating intermediate thoughts as tree nodes and reasoning outcomes as leaf nodes, iterating through node expansion and outcome simulation to derive optimized solutions (Zhang et al., 2024; Rabby et al., 2024; Zhang et al., 2025). Notably, Chang et al. (2025) incorporates self-refinement into node exploration, enabling MCTS to both identify the best node for subsequent iterations and optimize it for maximum exploitation.
5 Conclusion
We introduce AlphaApollo, a self-evolving agentic reasoning system that aims to tackle two core bottlenecks in foundation model reasoning—limited model-intrinsic capacity and unreliable, compute-intensive test-time iteration—by orchestrating multiple models with professional tools. By pairing exact computation (Python with scientific and symbolic libraries) with targeted retrieval, AlphaApollo produces verifiable, tool-augmented solutions, while a shared, evolving state enables parallel proposal, execution, testing, and refinement. Empirically, it delivers consistent improvements across AIME 2024/2025 and diverse model families. This release focuses on tool-augmented reasoning; upcoming versions will add multi-round, multi-model test-time scaling, and extensions to broader domains. We will open-source the full system and results to support reproducibility and accelerate community-driven progress toward trustworthy, real-world reasoning.
Appendix A Implementation Details of AlphaApollo
In this section, we provide the employed prompts in the AlphaApollo, including:
  • the system prompt of the main model (Figure 9);
  • the descriptions for the computational module (Figure 10) and the retrieval module (Figure 11);
  • the system prompts for the query rewriter (Figure 12) and result summarizer (Figure 13) in the retrieval module.
A.1 AlphaApollo System Prompt
System prompt of the main model You should solve the given math problem step by step.
There are some tools that you can use to help you solve the math problem.
Examine the available tools and determine which ones might relevant and useful for addressing the query. Make sure to consider the tool metadata for each tool.
Based on your thorough analysis, decide if your memory is complete and accurate enough to generate the final output, or if additional tool usage is necessary.
@@ A.1 AlphaApollo System Prompt
Figure 9: System prompt. In system prompt, we instruct the model to solve math problems step by step, inspect tool metadata to decide which tools to use, and present the final answer in box.
A.2 Tool Descriptions
Tool description of the computational module A tool to execute python code. You need to use print() to get the result of the code execution. If you are not familiar with the package (sympy, scipy, numpy, math, cmath, fractions, itertools) you are using, you can call ‘rag_system_retrieve’ tool before calling ‘python_code_executor’ tool.
Args:
    code: code text
Returns:
    The result of the code execution
metadata:
    “limitations”: 1) “Do not perform plotting operations, such as using matplotlib.”, 2) “If you want to use external packages, you can only use sympy, scipy, numpy, math, cmath, fractions, and itertools.”, 3) “Not applicable to geometry and number theory problems.”, 4) “No access to any system resources, file operations, or network requests.”, 5) “All calculations must be self-contained within a single function or script.”, 6) “Input must be provided directly in the query string.”, 7) “Output is limited to numerical results or simple lists/tuples of numbers.”.
    “best_practices”: 1) “Provide clear and specific queries that describe the desired mathematical calculation.”, 2) “Include all necessary numerical inputs directly in the query string.”, 3) “Ensure all required numerical data is included in the query.”.
Figure 10: Tool description of the computational module. In this tool description, we instruct the model on how to use computational module.
Tool description of the retrieval module A search tool system for querying Python package documentation.
Args:
    repo_name: One of [sympy, scipy, numpy, math, cmath, fractions, itertools]
    query: Natural-language query for retrieval, e.g., ‘Function interface and examples of calling sympy to solve nonlinear equations’
    top_k: Number of documents to return per sub-query (default: 3)
Returns:
    str: The query result, including relevant code examples, descriptions, and usage.
metadata:
    “limitations”: 1) “Not suitable for solving specific math problems.”
    “best_practices”: 1) “Ask for usage instructions or examples for a specific function interface whenever possible.”
Figure 11: Tool description of the retrieval module. In this tool description, we instruct the model on how to use retrieval module.
A.3 Prompts of the Retrieval Module
System prompt of the query rewriter You decide whether to generalize the user’s query. If it is already short, clear and general, return an empty line. Otherwise, rewrite it into ONE short, general question about the relevant functions/classes/APIs for this task type. The rewrite should:
(1) Keep only: task category and the library/tool name if present.
(2) Do NOT include any specific objects/structures, numbers, variable names, operators, or concrete equation forms.
(3) Return a single concise question text (no preamble), or empty if the original is already short, clear, and general.
Return ONLY the question text or empty.
Figure 12: System prompt of the query rewriter. In the query rewriter prompt, the model is asked to decide whether and how to generalize the user’s query.
System prompt of the result summarizer You decide whether to answer. Treat the provided docs as optional hints; they may be partial or off-topic. If the docs already clearly answer the query, return an empty line. Otherwise, write a concise, self-contained answer that:
(1) includes fully-qualified function or class names when possible (e.g., sympy.core.function.diff);
(2) includes brief, runnable usage examples;
(3) NEVER mentions files, tests, sources, or ’documents’, and avoids any meta commentary about what the docs do or do not contain;
(4) may rely on general knowledge beyond the docs to help the user.
Return ONLY the answer text (no preamble) or empty.
Figure 13: System prompt of the result summarizer. In the result summarizer prompt, the model is asked to decide whether and how to summarize the provided docs.
Appendix B Case Studies
B.1 Python Code Errors

### v2

@@ A.3 Test-time Iteration
Most early test-time scaling methods utilize parallel iteration, where the model generates multiple independent reasoning trajectories. The final answer is then selected through various aggregation mechanisms. Self-Consistency (wang2022self) employs simple voting, Semantic Self-Consistency (knappe2024semantic) leverages semantic similarity, CISC (taubenfeld2025confidence) utilizes confidence metrics, and MCR (yoran2023answering) implements model-decided final answers. Additionally, DIVSE (naik2023diversity) enhances the diversity of the reasoning trajectory by reformulating the original question to encourage exploration of a broader solution space. While this strategy enhances FM reasoning by allocating more computational budget to the reasoning process (snell2024scaling), it is fundamentally limited by the lack of interaction among different reasoning trajectories. This isolation limits the model in leveraging from previous attempts or refining its approach based on earlier outputs. The independence between trajectories constrains the potential for iterative improvement and fails to capitalize on insights that could emerge from comparing or combining intermediate reasoning steps (qi2025learning).
@@
Sequential iteration strategies leverage prior reasoning processes and outcomes to identify limitations and iteratively refine subsequent generations. Notably, Self-Refine methods (madaan2023self; qu2024recursive; chen2023teaching) prompt FMs to revise initial outputs, generating improved solutions. Similarly, muennighoff2025s1 employs a deliberation mechanism, replacing the end-of-sequence token with the keyword ‘wait’ to enable continuous reasoning post-answer generation. OPRO (yang2024large) iteratively optimizes prompts based on prior outcomes to yield superior solutions. While these approaches enhance reasoning, their reliance on self-correction without external supervision can lead to unreliable outcomes (huang2024large). In contrast, Reflexion (shinn2023reflexion), RCI (kim2023language), and Refiner (paul2023refiner) incorporate external feedback from environments, code executors, and critic models, respectively, providing robust supervision and improving performance. Furthermore, methods like Alpha-Evolve (novikov2025alphaevolve) and TextGrad (yuksekgonul2025optimizing) extend sequential iteration to scientific domains, demonstrating FMs’ potential in tackling complex challenges, such as code optimization and molecular synthesis.
@@
Mixed iteration strategies integrate parallel and sequential iteration to combine the exploratory breadth of independent trajectories with the refinement depth of conditioned generations. This hybrid approach optimizes test-time compute allocation, often surpassing pure parallel or sequential methods on complex reasoning tasks by balancing exploration (through parallel iteration) and exploitation (via sequential refinement). A prominent example is Monte Carlo Tree Search (MCTS) (silver2016mastering; silver2017mastering), which employs multi-round node expansion through parallel iteration and outcome simulation via sequential iteration. Recent advancements apply MCTS to enhance FM reasoning by treating intermediate thoughts as tree nodes and reasoning outcomes as leaf nodes, iterating through node expansion and outcome simulation to derive optimized solutions (zhang2024accessing; rabby2024mc; zhang2025llama). Notably, chang2025step incorporates self-refinement into node exploration, enabling MCTS to both identify the best node for subsequent iterations and optimize it for maximum exploitation.
Appendix B Agentic Reasoning: Implementation Details and Discussion
B.1 The Computational Module
Figure 11: The pipeline of the computational module in processing code.
  • SymPy (meurer2017sympy) is a computer algebra system for symbolic mathematics, enabling the manipulation and solution of mathematical expressions in closed form. For example, it can determine the exact roots of the cubic equation $x^{3}-2x+1$ within Python.
  • NumPy (numpy) is a fundamental library for fast, vectorized numerical computing, providing powerful support for $N$-dimensional arrays and matrices. Typical tasks include linear algebra operations such as computing the dot product of two matrices.
  • SciPy (scipy) is a comprehensive library of advanced numerical algorithms for science and engineering, supporting integration, ordinary differential equations, optimization, signal processing, sparse linear algebra, statistics, and more. For instance, it can readily solve problems like finding the minimum of $f(x)=\sin(x)+x^{2}$ over the interval [3,3][-3,3].
Python environment and code execution.
Error correction.
The model-generated code can contain errors. To improve the robustness against these errors, AlphaApollo’s computational module incorporates a rule-based error-correction. When the model generates Python code, the rule-based approach detects and automatically corrects errors identifiable through predefined rules, and then passes the corrected code to the Python interpreter for execution.
In rule-based error-correction, all IndentationError can be detected, with most being automatically correctable. The rule-based approach verifies the legality of indentation line by line, identifying and removing unnecessary space tokens to ensure proper indentation throughout the code. Additionally, certain SyntaxError, such as extraneous markdown blocks wrapping the code, can be detected and resolved. This method extracts the core Python code from wrapped content, ensuring the generated code is executable.
In addition, for errors such as ValueError, TypeError, AttributeError, and ImportError, which commonly arise when using external libraries (e.g., calling sympy.factorial with a negative argument, although it only accepts non-negative values), the computational module can invoke the retrieval module (Sec. B.2) for further guidance. By combining this capability with the two-fold error-correction mechanism, the computational module creates a model-friendly Python environment that empowers foundation models to perform code-augmented reasoning.
B.2 The Retrieval Module
  • Query rewriter. The workflow begins with the query rewriter, which transforms the initial query into a retrieval-friendly specification. Instead of passing detailed task descriptions directly, the rewriter abstracts them into generalizable forms that emphasize functional intent. For example, an over-detailed query “Solve the equation $12x^{2}-xy-6y^{2}=0$ for $x$ symbolically” is rewritten as a more suitable one, “Solve a quadratic equation symbolically for a variable.” This reformulation abstracts away numerical details while retaining the core intent, resulting in more relevant retrieval terms for the retriever. We implement the query rewriter using an instruction-following model with a tailored prompt template.
  • Document retriever. After rewriting the query, the document retriever searches for relevant information in the module’s indexed corpus, which includes the source code of a Python library and its associated documentation. Technically, we partition the corpus into overlapped chunks, which could improve retrieval effectiveness by aligning units with sentence-level semantics and reducing the inclusion of irrelevant content. We implement overlapped chunking with a fixed-length sliding window, allowing adjacent chunks to share overlapping tokens and thus preserve cross-boundary context. These chunks are then encoded by the embedding model and stored in a vector database. When the retriever receives a refined query from the rewriter, it first encodes the query using the same embedding model employed during document indexing. The retriever then conducts a cosine-similarity search over the database to locate the top-$K$ chunk embeddings relevant to the query. The corresponding document segments are retrieved and assembled into a composite context, which is then passed to the summarization module.
  • Result summarizer. Finally, the result summarizer filters and summarizes the retrieved context into a concise response. Its role is to highlight callable functions, required arguments, and minimal working examples rather than returning raw documentation. For instance, when the retriever provides (i) demonstrations of sympy.solve, (ii) module descriptions of solveset, and (iii) general API notes on sympy, the summarizer distills them into an actionable tool description. It identifies solve as the appropriate function, specifies its key arguments, and generates a compact example such as: from sympy import solve, symbols; a, b, c, x = symbols(’a b c x’); solve(ax**2 + bx + c, x).
B.3 Experimental Settings
In this setting, we enable tool-augmented reasoning as a baseline without any training or agentic evolution. The model performs multi-turn interactions with access to computational (Python code execution) and retrieval (local RAG) tools, limited to a maximum of 4 rounds ($T_{\text{max}}=4$) to facilitate step-by-step problem-solving. We set the temperature to 0.6 for balanced exploration, with history length capped at 8 to maintain context across turns. No additional learning or verification mechanisms are applied, and each question is evaluated in a single-pass manner.
B.4 Prompts
In this part, we provide the employed prompts in the AlphaApollo in Figures 13 and 14.
System prompt of the first turn You are a math problem solver agent tasked with solving the given math problem step-by-step.
Your question: {question}
@@ B.4 Prompts
Figure 13: In the system prompt of the first turn, we instruct the model to solve math problems step by step, decide which tools to use, and present the final answer in a boxed format.
System prompt for subsequent turns You are a math problem solver agent tasked with solving the given math problem step-by-step.
Your question: {question}
[…]
Figure 14: In the system prompt of subsequent turns, we integrate the memory from previous turns and instruct the model to continue solving the math problem step by step, decide which tools to use, and present the final answer in a boxed format.
Appendix C Agentic Learning: Implementation Details and Discussion
Prior agentic learning systems, such as VerlTool (jiang2025verltool) and RL-Factory (chai2025rlfactory), stabilize agentic learning by employing a masking strategy that assigns zero weight to tokens from tool responses, thereby excluding them from optimization. Formally, considering the trajectory sampling process of GRPO as an example, the objective for trajectory-level RL optimization is:
$\mathcal{J}_{\text{Trajectory-GRPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}p_{1}\sim\mathcal{D}\\ \{y^{(i)}_{k}\}_{i=1}^{N}\sim\pi_{\text{old}}(\cdot|p_{1})\end{subarray}}\Bigl[\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|y^{(i)}|}\sum_{k=1}^{|y^{(i)}|}\min\Big(m(y_{k}^{(i)})\frac{\pi_{\theta}\big(y_{k}^{(i)}\mid p_{1},y_{<k}^{(i)}\big)}{\pi_{\text{old}}\big(y_{k}^{(i)}\mid p_{1},y_{<k}^{(i)}\big)}A_{k}^{(i)},$ (4)
$\operatorname{clip}\big(m(y_{t}^{(i)})\frac{\pi_{\theta}\big(y_{k}^{(i)}\mid p_{1},y_{<k}^{(i)}\big)}{\pi_{\text{old}}\big(y_{k}^{(i)}\mid p_{1},y_{<k}^{(i)}\big)},1-\epsilon,1+\epsilon\big)A_{k}^{(i)}\Big)-\beta\mathbb{D}_{\mathrm{KL}}\big[\pi_{\theta}\big\|\pi_{\mathrm{ref}}\big]\Bigr]$
where $y_{k}^{(i)}$ denotes $k$-th token of the $(i)$-th sampled trajectory, $m(y_{k}^{(i)})=\mathds{1}\big(y_{k}^{(i)}\in\mathcal{O}\big)$ indicates the masking strategy that optimize the policy with the model-generation content $\mathcal{O}$.
Similarly, the SFT in trajectory-level optimization learns given the full agentic reasoning trajectory while masking tokens from external tools.
$\mathcal{J}_{\text{Trajectory-SFT}}(\theta)\;=\;\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{|y|}m(y_{k})\log\pi_{\theta}(y_{k}\mid y_{<k},x)\right],$ (5)
Although the masking strategy prevents the policy from learning tool responses, thereby enhancing stability, the long-horizon nature of agentic reasoning can still destabilize training. Low-probability tokens arising in extended trajectories may compromise optimization, leading to reward collapse, as evidenced in studies on agentic reinforcement learning (xue2025simpletir; wang2025ragen).
C.1 Experimental Settings
In this phase, we apply a learning rate of 1e-6, a training batch size of 128, and a group size of 8. The temperature is set to 0.4 to ensure stable performance monitoring throughout the reinforcement learning process. The training datasets consist of Lighteval-Math, LIMR, and DeepScaleR for comprehensive mathematical reasoning coverage. To demonstrate the generality of our framework, we trained on each of these datasets separately in our experiments.
C.2 Prompts
We use the same prompts as in Appendix B.4, shown in Figures 13 and 14.
Appendix D Agentic Evolution: Implementation Details and Discussion
D.1 Experimental Settings
Table 4: The detailed settings of models in agentic evolution. Here, output length refers to the maximum response length specified during evaluation of evolution.
xxModel Familyxx xxxxModel Namexx Maximum context length Output length
Qwen2.5 (yang2024qwen2) Qwen2.5-1.5B-Instruct 128K 8k
Qwen2.5-3B-Instruct 128K 8k
Qwen2.5-7B-Instruct 128K 8k
Qwen2.5-14B-Instruct 128K 8k
We show the configuration of agentic evolution in Table 4. This phase involves no training and utilizes a dual-model configuration. The policy model operates with a temperature of 0.7 to encourage exploration and diverse solution generation, while the verifier uses a temperature of 0.4. The self-evolving process runs for 10 rounds, the verifier conducts repeated verification for 5 times to ensure majority judgment. Both models share a maximum token limit of 8k to accommodate complex reasoning trajectories.
Prompt for Solver with $\mathcal{M}^{(r)}$ for the first turn You are a math problem solver agent tasked with solving the given math problem step-by-step.
Your question: {question}
Below are the previous solutions and their verification feedback: {previous_solutions}
[…]
Now it’s your turn to respond to the current step. You should first conduct the reasoning process. This process MUST be enclosed within <think> </think> tags. After completing your reasoning, choose only one of the following actions (do not perform multiple actions at the same time):
[…]
Figure 15: Prompt for Solver with $\mathcal{M}^{(r)}$ for the first turn. The system instruction for the first turn, incorporating the set of previous solutions and their validity ($\mathcal{M}^{(r)}$).
Prompt for Solver with $\mathcal{M}^{(r)}$ for subsequence turn You are a math problem solver agent tasked with solving the given math problem step-by-step.
Your question: {question}
Below are the previous solutions and their verification feedback: {previous_solutions}
[…]
Prior to this step, you have already taken {step_count} step(s). Below is the interaction history: {memory_context}
Now it’s your turn to respond to the current step. You should first conduct the reasoning process. This process MUST be enclosed within <think> </think> tags. After completing your reasoning, choose only one of the following actions (do not perform multiple actions at the same time):
[…]
Figure 16: Prompt for Solver with $\mathcal{M}^{(r)}$ for the subsequent turns. Instructions for subsequent turns where the agent must integrate the interaction history (memory_context) with the evolutionary feedback from $\mathcal{M}^{(r)}$ to advance the solution step-by-step.
Prompt for Abstractor ROLE
You are a Mathematical Logic Auditor. Compress the interaction log into a Verification Brief.
INPUT DATA
The log contains ‘<think>‘, ‘<python_code>‘, ‘<tool_response>‘, and ‘<answer>‘ tags.
PROTOCOL
1. Filter Noise: Retain only the mathematical setup, derived constants, and successful logic. Discard syntax errors, backtracking, and internal monologue.
2. Track Origins: Explicitly differentiate between values computed via code and those asserted via intuition or external knowledge.
3. Format: Use LaTeX for math. Present as sequential Logical Checkpoints.
OUTPUT TEMPLATE
Strategy: [Brief summary of the approach]
Logical Checkpoints:
1. Setup: [Variable definitions and initial conditions]
2. Intermediate Result: [Key derived values]
3. Pivotal Step: [Crucial logic or calculation]
4. Resolution: [How the final result was reached]
Final Claim: <answer>…</answer>
TASK
Summarize:
{content}
Figure 17: System instructions for Abstractor. The agent generates a structured summary by extracting the key mathematical setup, intermediate results, and the pivotal logic leading to the final answer, while discarding irrelevant content.
Prompt for Evaluator for the first turns. You are a math verifier agent whose only job is to check whether the solver agent’s proposed solution is correct.
Original question:
{question}
Solver’s latest solution attempt:
{solver_solution}
Now it’s your turn to respond to the current step. You should first conduct the reasoning process. This process MUST be enclosed within <think></think>tags. After completing your reasoning, choose only one of the following actions (do not perform both):
1) <python_code>…</python_code>: If computation/checking is helpful, emit exactly ONE <python_code>…</python_code>block with pure Python 3. Inspect the <tool_response>(stdout from your code). If it disagrees with your reasoning, correct yourself.
2) <local_rag>…</local_rag>: You have access to a RAG System tool to search for documentation or examples (Supported repos: sympy, scipy, numpy, math, cmath, fractions, itertools). Emit exactly ONE <local_rag>…</local_rag>block with a JSON object. Inspect the returned <tool_response> (RAG result). If it disagrees with your reasoning, correct yourself. For example: <local_rag>{{”repo_name”: ”sympy”, ”query”: ”your query here”}}</local_rag>.
3) <report>…</report>: If you are ready to conclude, wrap your verification report inside <report>…</report>tags. The report should:
- Clearly state whether the policy solution appears correct or incorrect.
- Explain the key reasoning behind your judgment (keep it concise).
[…]
- The judgement should be enclosed within <report>…</report>tags.
Figure 18: The system instructions for the evaluator agent for the first turn. It receives the problem and the candidate solution, utilizes available tools to audit the logic, and outputs a verification.
Prompt for Evaluator for the subsequence turns. You are a math verifier agent whose only job is to check whether the solver agent’s proposed solution is correct.
Original question:
{question}
Solver’s latest solution attempt:
{solver_solution}
Prior to this step, you have already taken {step_count} step(s). Below is the interaction history: {memory_context}
Now it’s your turn to respond to the current step. You should first conduct the reasoning process. This process MUST be enclosed within <think></think>tags. After completing your reasoning, choose only one of the following actions (do not perform both):
1) <python_code>…</python_code>: If computation/checking is helpful, emit exactly ONE <python_code>…</python_code>block with pure Python 3. Inspect the <tool_response>(stdout from your code). If it disagrees with your reasoning, correct yourself.
2) <local_rag>…</local_rag>: You have access to a RAG System tool to search for documentation or examples (Supported repos: sympy, scipy, numpy, math, cmath, fractions, itertools). Emit exactly ONE <local_rag>…</local_rag>block with a JSON object. Inspect the returned <tool_response> (RAG result). If it disagrees with your reasoning, correct yourself. For example: <local_rag>{{”repo_name”: ”sympy”, ”query”: ”your query here”}}</local_rag>.
3) <report>…</report>: If you are ready to conclude, wrap your verification report inside <report>…</report>tags. The report should:
- Clearly state whether the policy solution appears correct or incorrect.
- Explain the key reasoning behind your judgment (keep it concise).
[…]
- The judgement should be enclosed within <report>…</report>tags.
Figure 19: The system instruction for Evaluator for subsequent turns.
Prompt for Summarizer You are a math verification report aggregator. Multiple evaluators have independently verified a solver agent’s solution and reached the same judgment. Your task is to synthesize their reports into a single, comprehensive report.
Original question:
{question}
Solver agent’s solution:
{Solver_solution}
[…]
Below are the individual reports:
{individual_reports}
Your task:
1. Analyze the key reasoning points from each verifier report.
2. Synthesize the most compelling arguments and evidence into a single, coherent report.
3. Ensure the aggregated report is comprehensive yet concise.
4. Maintain the same final judgment as the individual reports.
Wrap your aggregated report inside <report>…</report>tags. The report should:
- Clearly state whether the policy solution is correct or incorrect.
- Combine the strongest reasoning from all verifier reports.
[…]
Figure 20: Instructions for summarizing verifications. The agent summarizes the key information from several concordant evaluators into a single representative judgment, maintaining the original consensus judgment.
D.2 Prompts
In this part, we present the detailed system prompts governing the specialized agents within our agentic evolution framework (Sec. 2.3). These prompts dictate the specific behaviors, tool permissions, and interaction protocols for each role.
Solver: The Solver is responsible for generating solution trajectories. In the initial round, the prompts follow the standard protocols shown in Figures 13 and 14 (Sec. B.4). In subsequent rounds, the solver is augmented with long-term memory entities. Figure 15 illustrates the initialization prompt containing the history of past attempts ($\mathcal{M}^{(r)}$), while Figure 16 shows the instruction for intermediate steps, which integrates both the previous attempts and the local interaction context.
Abstractor: To facilitate efficient memory storage, the Abstractor (Figure 17) is tasked with compressing raw interaction logs. It filters out computational noise, syntax errors, and backtracking, restructuring the successful reasoning path into concise “logical checkpoints”. This process distinguishes between computed values and intuitive assertions, ensuring the stored memory is high-quality and retrieval-ready.
Evaluator: As shown in Figure 18, the agent is instructed to verify the solver’s candidate solution against the original problem. The prompt enforces a ”Think-Check-Report” lifecycle, explicitly authorizing the use of tools to validate intermediate steps and definitions before issuing a binary correctness judgment. For multi-turn verification sessions, the prompt in Figure 19 adapts to include the interaction history.
Summarizer: Designed for ensemble verification scenarios, the Summarizer (Figure 20) is activated when multiple evaluators reach a consensus. It aggregates individual verification into a single judgement by synthesizing the arguments and removing redundancy, thereby providing a comprehensive justification.
Appendix E Additional Experimental Results
In this section, we provide additional results of the experiments.
E.1 Agentic evolution
Table 5: Agentic evolution results (accuracy in %). w/o Evo uses tools without evolution; +Evo enables agentic evolution with tools. Bold marks the better result.
Datasets Qwen3-4B-Instruct-2507
w/o Evo +Evo
AIME24 44.98 81.68
AIME25 32.50 61.68
CMIMC 25 16.25 16.25
HMMT25 Feb 15.83 43.33
HMMT25 Nov 23.30 42.53
BRUMO25 24.15 51.68
SMT 2025 30.65 32.58
Average 26.81 47.10
[…]
Additional evidence on the Qwen3 series confirms the benefit of agentic evolution.
Additionally, we conduct experiments on Qwen3-4B-Instruct (Table 5) to further validate the effect of evolution. Enabling evolution (+Evo) substantially outperforms the non-evolving setting (w/o Evo), improving average accuracy from 26.81% to 47.10% (+20.29). The largest gains appear on AIME24 (+36.70+36.70) and AIME25 (+29.18+29.18), with similarly strong improvements on HMMT25 Feb (+27.50+27.50), BRUMO25 (+27.53+27.53), and HMMT25 Nov (+19.23+19.23). Overall, these results on the Qwen3 series further demonstrate that agentic evolution is robust and consistently improves mathematical reasoning with tools.
Appendix F Case Studies
F.1 Python Code Errors